OpenStudy (anonymous):

Find the domain and range of a real function f(x) = (x^2-1)/(x+1)?

OpenStudy (lgbasallote):

you have \[\frac{x^2 - 1}{x+1}\] the domain restriction will be the value of x that will make x + 1 = 0

OpenStudy (mathteacher1729):

Domain: zero in denominator? log of negative number or zero? sqrt of a negative number? Range: Graphing or Finding domain of inverse or arguments from the form of the function itself (a bit more advanced)

OpenStudy (anonymous):

Domain = R-[-1]

OpenStudy (anonymous):

Range=??

OpenStudy (anonymous):

see x!=-1 that's domain and range is x!=-2

OpenStudy (mathteacher1729):

The range is all values of y that the function can take on. \[y = \frac{(x-1)(x+1)}{(x+1)}\] we can cancel and write y = x - 1 This has range of all real numbers. BUT our original function did not allow for x = -1 which means y cannot equal (-1 - 1) = -2. So the range is all real numbers except y = -2.

OpenStudy (anonymous):

but the answer in my book says range = R

OpenStudy (mathteacher1729):

That is incorrect. If the range was all real number , that means we can solve for ANY real number we choose. Solve \[\huge -2 = \frac{x^2-1}{x+1}\] You cannot do this. You would need to have x = -1 after canceling common factors, and ... x cannot equal -1.

OpenStudy (mathteacher1729):

See attached.

OpenStudy (anonymous):

I m on iPad now and this is not allowin me to draw or write equation it is also getting unstable sorry can't help write now

OpenStudy (anonymous):

@satellite73 plzz help

OpenStudy (anonymous):

i am confused plzz help

OpenStudy (anonymous):

the answer in the book is wrong as mathteacher said pity

OpenStudy (anonymous):

Can u give the explanation

OpenStudy (anonymous):

is the Domain = R-[-1]???

OpenStudy (anonymous):

Agreed even I think book is wrong

OpenStudy (anonymous):

\[f(x)=\frac{x^2-1}{x+1}=\frac{(x+1)(x-1)}{x+1}=x-1\] only provided \(x\neq -1\)

OpenStudy (anonymous):

then R=??

OpenStudy (anonymous):

if \(x=-1\) you would get \[f(-1)=\frac{(-1)^2-1}{-1+1}=\frac{1-1}{-1+1}=\frac{0}{0}\] and although you can write it with pencil and paper, \(\frac{0}{0}\) is not a number that is why the domain excludes \(-1\) because you cannot divide by zero

OpenStudy (anonymous):

ok got it then range=?

OpenStudy (anonymous):

therefore the domain is "all real numbers except -1" which can be written as \[\mathbb{R}-[-1]\] if you like

OpenStudy (anonymous):

Then Range=?

OpenStudy (anonymous):

now for the range \(f(x)=x-1\) for all \(x\) except \(-1\) where the function is undefined \(y=x-1\) is a line with slope 1 and \(y\) intercept \(-1\) which is exactly what the graph looks like therefore the range is all real numbers, except the number you would get if you replace \(x\) by \(-1\) because \(-1\) is not in the domain if you replace \(x\) by \(-1\) you get \[f(-1)=-1-1=-2\] and so this function can never be \(-2\) because you may not replace \(x\) by \(-1\)

OpenStudy (anonymous):

so Range= R-[-2]

OpenStudy (anonymous):

another way to say this is that there is no input for the function \[f(x)=\frac{x^2-1}{x+1}\] that would give the answer \(-2\) if you want we could try and solve it, and see why it doesn't work yes, it is what you wrote

OpenStudy (anonymous):

thxxx

OpenStudy (anonymous):

yw hope it is clear

OpenStudy (anonymous):

Sorry cudnt help yahoo